OpenStudy (anonymous):

solve the system of equations: 4x-8y =0 & x^2 + y^2 = 25

OpenStudy (bradely):

4x-8y =0 4x=8y x=2y x^2 + y^2 = 25 (2y)^2+y^2 = 25 5y^2=25 y^2=5 y=+ or- sqrt(5) x=+ or -2 sqrt(5) Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

thanks!